OpenStudy (anonymous):

Ian solved the system of equations by making a substitution for y in the second equation. What equation did he get as a result of that substitution? y = –3x + 1 5x – 2y = 31

OpenStudy (anonymous):

5x + 6x + 1 = 31 5x + 6x – 2 = 31 –15x + 5x = 31 –15x – x = 31

OpenStudy (anonymous):

B??

OpenStudy (butterfield1215):

i will give you a site that you can use that will help you with this kin of stuff its called mathway.com and it checks your work and i will check to see if your right

OpenStudy (butterfield1215):

your correct

OpenStudy (anonymous):

5x – 2y = 31 5x - 2(–3x + 1) = 31 5x +6x -2 = 31

OpenStudy (anonymous):

Off course you guessed it right :)

OpenStudy (anonymous):

Thank you!